# Optimizing Parameters for Static Equilibrium of Discrete Elastic Rods with Active-Set Cholesky

### v1

@@ Abstract
We propose a parameter optimization method for achieving static equilibrium of discrete elastic rods. Our method simultaneously optimizes material stiffness and rest shape parameters under box constraints to exactly enforce zero net force while avoiding stability issues and violations of physical laws. For efficiency, we split our constrained optimization problem into primal and dual subproblems via the augmented Lagrangian method, while handling the dual subproblem via simple vector updates. To efficiently solve the box-constrained primal subproblem, we propose a new active-set Cholesky preconditioner. Our method surpasses prior work in generality, robustness, and speed.
Index Terms:
Parameter optimization, hair simulation, inverse problem, active-set method
@@ I Introduction
When modeling deformable objects or fabricating elastic materials, sagging under gravity or external loads is a well-known problem. This behavior arises because gravity is often implicitly considered during modeling but simulators typically neglect it during initialization [1, 2, 3, 4, 5]. While stiffening materials can reduce such sagging, it alters the dynamic response of the elastic material. Alternatively, modifying rest shape parameters can achieve static equilibrium although substantial rest shape changes can introduce stability problems or increase the likelihood of undesirable rest configurations.
@@
To address the sagging problem, we propose a new parameter optimization method that is guaranteed (at solver convergence) to achieve static equilibrium of DER-based strands. Our method simultaneously optimizes rest shape and material stiffness parameters while minimizing parameter changes and respecting their box constraints. We formulate our problem as a constrained minimization and decompose it into primal and dual subproblems via the augmented Lagrangian method (ALM) [15], addressing the primal subproblem with a Newton-type optimizer and the dual subproblem via vector updates. To efficiently and accurately handle the primal subproblem, we take the box constraints into account and solve the symmetric positive definite (SPD) inner systems using a variant of Conjugate Gradient (CG) with active sets, i.e., Modified Proportioning with Reduced Gradient Projections (MPRGP) [16]. Additionally, we propose a new preconditioner based on a Cholesky-type solver with active sets to accelerate the convergence of MPRGP. Figure 1 shows our method in action.
Figure 1: Our method optimizes material stiffness and rest shape parameters to achieve static equilibrium for DER-based strands. It preserves the original hairstyle rather than sagging, demonstrates natural dynamic strand behaviors in response to prescribed motions of the root vertices, and eventually restores the strands to their initial configuration. Our parameter optimization required 713 seconds for 3.2k strands; forward simulation without collision handling took 25 seconds per frame.
@@ II-A Elastic Rod Simulation
To capture bending and twisting of one-dimensional elastic strand structures, Pai [17] introduced the Cosserat theory into graphics, which was later extended to dynamical systems [18]. The Cosserat theory has since been adopted within position-based dynamics [19] and further extended with quaternions [20, 4], volume constraints [21], and projective dynamics [22]. Instead of evolving edge rotations, Shi et al. [23] proposed using a volume-like torsion energy to capture twisting effects. Curvature-based approaches were also presented to capture elastic rod dynamics with a smaller number of degrees of freedom (DOFs) [24, 25]. From the various methods for simulating elastic strands, we selected DER [6, 7] because it can handle general bending and twisting cases with few DOFs.
@@ II-B Sag-Free Simulation
To achieve sag-free simulations (or equivalently, static equilibrium), one common approach is to enforce zero net forces in the dynamical system. Hadap [26] proposed using inverse dynamics to solve nonlinear force equations for reduced multibody systems, although such methods may not necessarily be applicable to maximal coordinate systems. To solve nonlinear force equations, the Asymptotic Numerical Method (ANM) was employed, achieving faster convergence compared to Newton-type optimizers [27, 28], although it remains unclear how to incorporate box constraints, which help prevent stability problems and violations of physical laws. Curvature-based solvers have also been employed, since they need fewer DOFs to represent elastic rod dynamics. Derouet-Jourdan et al. [29] presented a method for achieving static equilibrium of curvature-based elastic strands [24] and extended their work with frictional contacts [30]. The curvature-based formulations have extensively been used for inverse problems to model, design, and fabricate elastic rods [31, 32, 33, 34].
@@
To achieve static equilibrium in more general contexts, Twigg and Kačić-Alesić proposed minimizing the force norm by optimizing rest shape parameters in mass-spring systems consisting of one-dimensional elements [1]. Takahashi and Batty extended their approach to DER [5], in work closely related to ours. We clarify key differences in Sec II-C. Rest shape optimization has also been applied to two-dimensional sheet-like structures [35] and three-dimensional deformable volumes [36, 37]. To enforce zero-net-force constraints, adjoint methods have been extensively used [2, 38, 12]. While these methods are efficient, their implementation tends to be significantly more complicated because they involve Hessians of objective terms due to differentiation of the nonlinear force constraints. Given that DER Hessian treatments are the subject of ongoing study [12, 23], we formulate our problem without explicitly incorporating the Hessian into the parameter optimization. While finite differences could bypass this complexity [39], they are known to be significantly slower than optimization with analytical gradients.
@@
As an alternative, Hsu et al. [3, 4] proposed a global-local initialization method that first computes global forces to achieve static equilibrium and then adjusts local elements to generate such forces. However, this approach would require excessive stiffening of materials and is applicable only to simulation approaches where local force elements can exclusively be defined. As such, the global-local initialization is not applicable to DER, as discussed by Takahashi and Batty [5].
II-C Differences from [5]
@@
There are several key differences between our work and the closely related prior work by Takahashi and Batty [5]. Their approach optimizes only the rest shape parameters and may fail to achieve zero net forces. By contrast, our method simultaneously optimizes both the material stiffness and rest shape parameters while enforcing zero net forces as a hard constraint, ensuring static equilibrium at solver convergence. We also reduce the computational cost of parameter optimization in two ways. First, we order the optimization parameters in an interleaved manner to yield a banded system, leveraging the one-dimensional structure of strands, which eliminates the need for matrix reordering in Cholesky solvers. Second, by considering the overlapping force spaces with 4D curvatures [6, 9], we include only two rest curvature variables per vertex, reducing the DOF count in the optimization (see Sec. IV for details). Moreover, instead of relying on penalty-based approaches for box constraints [5], we utilize an active-set-based iterative solver, MPRGP [16], to enforce box constraints precisely. We also propose a new active-set Cholesky preconditioner to accelerate the convergence of MPRGP.
III Discrete Elastic Rods Preliminary
@@
For completeness, we briefly summarize key details of the DER formulation [6, 7] before explaining our parameter optimization method (Sec. IV) and active-set Cholesky preconditioner (Sec. V-A). In our framework, we exclude contacts and thus process each strand independently, focusing below on a single strand.
@@
Given a strand discretized into $N$ vertices with positions $\mathbf{x}=[\mathbf{x}_{0}^{T},\ldots,\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{3N}$ and $N-1$ edges with angles $\bm{\theta}=[\bm{\theta}_{0},\ldots,\bm{\theta}_{N-2}]^{T}\in\mathbb{R}^{N-1}$, the generalized positions can be defined by interleaving vertex and edge variables as $\mathbf{q}=[\mathbf{x}_{0}^{T},\bm{\theta}_{0},\ldots,\bm{\theta}_{N-2},\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{4N-1}$. This arrangement leads to a banded Hessian due to locally defined DER objectives and time-parallel transport [7]. For simplicity, we assume a strand with a perfectly circular cross-section and uniform radius $r$. The root end of the strand is minimally clamped, fixing $\mathbf{x}_{0},\bm{\theta}_{0}$, and $\mathbf{x}_{1}$, which yields $84N-$ 8 active DOFs [5].
@@
We define a minimization problem for a forward simulation step with a DER objective $E(\mathbf{q})$ [5] as $\mathbf{q}=\operatorname*{arg\,min}_{\mathbf{q}}E(\mathbf{q})$, where $E(\mathbf{q})=E_{\mathrm{in}}(\mathbf{q})+E_{\mathrm{st}}(\mathbf{q})+E_{\mathrm{be}}(\mathbf{q})+E_{\mathrm{tw}}(\mathbf{q})$, and $E_{\mathrm{in}}(\mathbf{q})$, $E_{\mathrm{st}}(\mathbf{q})$, $E_{\mathrm{be}}(\mathbf{q})$, and $E_{\mathrm{tw}}(\mathbf{q})$ represent inertia, stretching, bending, and twisting objectives, respectively.
III-A Inertia
We define the inertia objective and its gradient as follows:
@@ III-A Inertia
$E_{\mathrm{in}}(\mathbf{q})=\frac{1}{2\Delta t^{2}}\left\lVert\mathbf{q}-\mathbf{q}^{*}\right\rVert^{2}_{\mathbf{M}},\nabla E_{\mathrm{in}}(\mathbf{q})=\frac{\mathbf{M}}{\Delta t^{2}}(\mathbf{q}-\mathbf{q}^{*}),$ (1)
@@
where $\mathbf{q}^{*}=\mathbf{q}^{t}+\Delta t\mathbf{\dot{q}}^{t}+\Delta t^{2}\mathbf{M}^{-1}\mathbf{f}_{\mathrm{ext}}$, $\mathbf{q}^{t}$ and $\mathbf{\dot{q}}^{t}$ denote the generalized positions and velocities at time $t$, respectively, $\Delta t$ time step size, $\mathbf{M}$ a diagonal generalized mass matrix [7, 40], and $\mathbf{f}_{\mathrm{ext}}$ generalized external forces.
III-B Stretching
We define the stretching energy [7] as $E_{\mathrm{st}}(\mathbf{q})=\sum_{i=1}^{N-2}E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})$, where $E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})$ is given by
$E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})=\frac{1}{2}\left(\frac{s\bm{\alpha}_{i}\pi r^{2}}{\mathbf{\bar{l}}_{i}}\right)(\mathbf{l}_{i}-\mathbf{\bar{l}}_{i})^{2},$ (2)
@@ III-B Stretching
with $s$ denoting a global constant scalar that scales stiffness parameters, $\bm{\alpha}_{i}$ as the stiffness parameter for stretching on edge $i$, $\mathbf{l}_{i}=\left\lVert\mathbf{x}_{i+1}-\mathbf{x}_{i}\right\rVert_{2}$ representing the length of edge $i$, and $\mathbf{\bar{l}}_{i}$ as its rest length. The gradient can be computed as
$\nabla_{\mathbf{x}_{i+1}}E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})$ $=-\nabla_{\mathbf{x}_{i}}E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})$ (3)
$=s\bm{\alpha}_{i}\pi r^{2}(\mathbf{l}_{i}\mathbf{\bar{l}}_{i}^{-1}-1)\mathbf{t}_{i},$ (4)
where $\mathbf{t}_{i}=\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i}}{\mathbf{l}_{i}}$ denotes the unit tangent vector of edge $i$.
III-C Bending
@@
$E_{\mathrm{be},i}(\mathbf{y}_{i})=\frac{1}{2}\left(\frac{s\bm{\beta}_{i}\pi r^{4}}{4(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)\left\lVert\bm{\kappa}_{i}-\bm{\bar{\kappa}}_{i}\right\rVert_{2}^{2},$ (5)
and $\mathbf{y}_{i}=(\mathbf{x}_{i-1}^{T},\bm{\theta}_{i-1},\mathbf{x}_{i}^{T},\bm{\theta}_{i},\mathbf{x}_{i+1}^{T})^{T}\in\mathbb{R}^{11}$, $\bm{\beta}_{i}$ denotes the bending stiffness at vertex $i$, and $\bm{\kappa}_{i}$ and $\bm{\bar{\kappa}}_{i}$ are the four-dimensional curvature and rest curvature, respectively [6, 9]. (While two-dimensional curvatures and rest curvatures can be employed with averaged material frames [7], this method has been shown to inadequately evaluate strand bending due to non-unit averaged material frames [41, 12].) The 11-dimensional gradient of (5) is given by [5]
$\nabla E_{\mathrm{be},i}(\mathbf{y}_{i})=\left(\frac{s\bm{\beta}_{i}\pi r^{4}}{4(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)\mathbf{J}_{\mathrm{cu},i}^{T}(\bm{\kappa}_{i}-\bm{\bar{\kappa}}_{i}),$ (6)
where $\mathbf{J}_{\mathrm{cu},i}\in\mathbb{R}^{4\times 11}$ denotes the Jacobian of $\bm{\kappa}_{i}$ with respect to $\mathbf{y}_{i}$. Here, $\mathbf{J}_{\mathrm{cu},i}^{T}(\bm{\kappa}_{i}-\bm{\bar{\kappa}}_{i})=\sum_{j=0}^{3}(\bm{\kappa}_{i,j}-\bm{\bar{\kappa}}_{i,j})\nabla\bm{\kappa}_{i,j}$, where $\bm{\kappa}_{i,j}$ denotes the $j$th entry of $\bm{\kappa}_{i}$.
III-D Twisting

### v2

@@ Abstract
We propose a parameter optimization method for achieving static equilibrium of discrete elastic rods. Our method simultaneously optimizes material stiffness and rest shape parameters under box constraints to exactly enforce zero net forces while avoiding stability issues and violations of physical laws. For efficiency, we split our constrained optimization problem into primal and dual subproblems via the augmented Lagrangian method, while handling the dual maximization subproblem via simple vector updates. To efficiently solve the box-constrained primal minimization subproblem, we propose a new active-set Cholesky preconditioner for variants of conjugate gradient solvers with active sets. Our method surpasses prior work in generality, robustness, and speed.
@@ I Introduction
When modeling deformable objects or fabricating elastic materials, sagging under external forces, such as gravity and applied loads, can ruin specific shapes carefully designed by artists and designers or captured from real world counterparts. This sagging problem arises because the external forces are often implicitly considered during modeling but simulators typically neglect it during initialization [1, 2, 3, 4, 5]. To prevent this issue, we aim to achieve static equilibrium of the objects preserving their original shapes at initialization. While stiffening materials can reduce the sagging, it alters the dynamic response of the elastic material. Alternatively, modifying rest shape parameters can achieve static equilibrium, though substantial rest shape changes can introduce stability problems or increase the likelihood of undesirable rest configurations [5].
@@
To address the sagging problem, we propose a new parameter optimization method that is guaranteed (at solver convergence) to achieve static equilibrium of DER-based strands. Our method simultaneously optimizes rest shape and material stiffness parameters while minimizing changes in the parameters and satisfying their corresponding box constraints; the latter are imposed to avoid violating physical laws or introducing stability issues in forward simulation. We formulate our problem as a constrained minimization and decompose it into primal and dual subproblems via the augmented Lagrangian method (ALM) [15], addressing the primal minimization subproblem with a Newton-type optimizer and the dual maximization subproblem via vector updates. To efficiently and accurately handle the primal subproblem, we take the box constraints into account and solve the symmetric positive definite (SPD) inner systems using a variant of conjugate gradient (CG) with active sets, modified proportioning with reduced gradient projections (MPRGP) [16]. Additionally, we propose a new preconditioner based on a Cholesky-based direct solver with active sets to accelerate the convergence of MPRGP. Figure 1 shows our method in action.
Figure 1: Our method optimizes material stiffness and rest shape parameters to achieve static equilibrium for DER-based strands. It preserves the original hairstyle without sagging, demonstrates natural dynamic strand behaviors in response to prescribed motions of the root vertices, and eventually restores the strands to their initial configuration. Our parameter optimization required 713 seconds for 3.2k strands; forward simulation without collision handling took 25 seconds per frame.
@@ II-A Elastic Rod Simulation
To capture bending and twisting of one-dimensional elastic strand structures, Pai [17] introduced the Cosserat theory into graphics, which was later extended to dynamical systems [18]. The Cosserat theory has since been adopted within position-based dynamics [19] and further extended with quaternions [20, 4], volume constraints [21], and projective dynamics [22]. Instead of evolving edge rotations, Shi et al. [23] proposed using a volume-like torsion energy to capture twisting effects. Curvature-based approaches were also presented to capture elastic rod dynamics with a smaller number of degrees of freedom (DOF) [24, 25]. From the various methods for simulating elastic strands, we selected DER [6, 7] because it can handle general bending and twisting cases with few DOFs.
@@ II-B Sag-Free Simulation
To achieve sag-free simulations (or equivalently, static equilibrium), one common approach is to enforce zero net forces in the dynamical system. Hadap [26] proposed using inverse dynamics to solve nonlinear force equations for reduced multibody systems, although such methods may not necessarily be applicable to general maximal coordinate systems (e.g., finite-element-method-based deformables and DER). To solve nonlinear force equations, the asymptotic numerical method (ANM) was employed, achieving faster convergence compared to Newton-type optimizers [27, 28], although it remains unclear how to incorporate box constraints, which help prevent stability problems and violations of physical laws. Curvature-based solvers have also been employed, since they need fewer DOFs to represent elastic rod dynamics. Derouet-Jourdan et al. [29] presented a method for achieving static equilibrium of curvature-based elastic strands [24] and extended their work with frictional contacts [30]. The curvature-based formulations have extensively been used for inverse problems to model, design, and fabricate elastic rods [31, 32, 33, 34].
@@
To achieve static equilibrium in more general contexts, Twigg and Kačić-Alesić proposed minimizing the force norm by optimizing rest shape parameters in mass-spring systems consisting of one-dimensional elements [1]. Takahashi and Batty extended their approach to DER [5] and proposed optimizing the rest length, rest curvature, and rest twist parameters while penalizing the force norm in the kinetic energy metric. However, this approach would fail to achieve static equilibrium because fully optimized rest shape parameters alone could still be insufficient to achieve zero net forces. Sag-free DER simulations have been attempted in other previous work as well [35, 10], albeit without offering formulation and algorithmic specifics. Rest shape optimization has also been applied to two-dimensional sheet-like structures [36] and three-dimensional deformable volumes [37, 38]. To enforce zero-net-force constraints, adjoint methods have been extensively used [2, 39, 12]. While this approach can be efficient and can ensure stable local minima by examining the Hessian of the DER objectives, handling the full Hessian due to differentiation of nonlinear force constraints can be complicated [12, 9]. Given that DER Hessian treatments are the subject of ongoing study [12, 23], we formulate our problem without explicitly incorporating the Hessian into the parameter optimization. While finite differences could bypass this complexity [40], they are known to be significantly slower than optimization with analytical gradients [15].
@@
As an alternative, Hsu et al. [3, 4] proposed a global-local initialization method that first computes global forces to achieve static equilibrium and then adjusts local elements to generate such forces. However, this approach would require excessive stiffening of materials and is applicable only to specific simulation approaches that can locally define forces without coupling neighboring elements. As such, the global-local initialization is not applicable to DER [5].
II-C Preconditioning with Active Sets
Preconditioners have been extensively used to reduce the number of necessary iterations for iterative solvers, such as CG. Preconditioners are also applicable and effective for variants of CG with an active-set method, e.g., MPRGP (which updates active sets in each iteration to find their optimal sets and solution) [16], by taking active sets into account during the preconditioning.
While Narain et al. [41] used modified incomplete Cholesky (MIC) preconditioning [42] for MPRGP, the derivations and rationale behind their approach are not explained, and it is not necessarily clear from their publicly available source code how to adapt it to support variants of Cholesky solvers (e.g., LDLT-type decomposition with $2\times 2$ diagonal blocks [43]). Recently, Takahashi and Batty [44] proposed a smoothed aggregation algebraic multigrid (SAAMG) preconditioner for MPRGP, using an SPAI-0 smoother [45] with active sets (which can easily be modified to a diagonal, weighted Jacobi, or symmetric Gauss-Seidel (SGS) smoother/preconditioner). However, their focus is on systems with M-matrices, and thus the effectiveness of their preconditioning on non-M-matrices, such as those in our formulation, remains unverified.
III Contributions
@@
We propose optimizing both the material stiffness and rest shape parameters simultaneously while enforcing zero net forces as a hard constraint via ALM to ensure the static equilibrium of DER-based strands at solver convergence. To reduce the optimization cost, we propose two techniques. First, we arrange the optimization parameters in an interleaved way to yield a banded system by leveraging the one-dimensional strand structure. This arrangement eliminates the need for matrix reordering in Cholesky-based solvers. Second, given the overlapping force spaces with 4D curvatures in the bending formulation [6, 9], we consider only two rest curvature variables per vertex in the optimization, reducing the DOF count and improving efficiency. Moreover, instead of using penalty-based methods [5], we employ an active-set-based iterative solver, MPRGP [16], to enforce the box constraints precisely. We also propose a new active-set Cholesky preconditioner to accelerate the convergence of MPRGP. Because of these improvements, our method can more robustly and efficiently achieve static equilibrium of DER-based strands.
IV Discrete Elastic Rods Preliminaries
@@
For completeness, we briefly summarize key details of the DER formulation [6, 7] before explaining our parameter optimization method (Sec. V) and active-set Cholesky preconditioner (Sec. VI-A). In our framework, we exclude contacts (as our goal is to achieve static equilibrium which introduces no contacts if none exist at initialization) and thus process each strand independently, focusing on a single strand below.
@@
Given a strand discretized into $N$ vertices with positions $\mathbf{x}=[\mathbf{x}_{0}^{T},\ldots,\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{3N}$ and N1N-1 edges with angles $\boldsymbol{\theta}=[\theta_{0},\ldots,\theta_{N-2}]^{T}\in\mathbb{R}^{N-1}$, the generalized positions can be defined by interleaving vertex and edge variables as $\mathbf{q}=[\mathbf{x}_{0}^{T},\theta_{0},\ldots,\theta_{N-2},\mathbf{x}_{N-1}^{T}]^{T}\in\mathbb{R}^{4N-1}$. This arrangement leads to a banded Hessian due to the locally defined DER objectives and time-parallel transport [7]. For simplicity, we assume a strand with a perfectly circular cross-section and uniform radius $r$. The root end of the strand is minimally clamped, fixing $\mathbf{x}_{0},\boldsymbol{\theta}_{0}$, and $\mathbf{x}_{1}$, which yields 4N84N-8 active DOFs [5]. For more details, we refer to a book on DER [46].
@@
We define a minimization problem for a forward simulation step with a DER objective $E(\mathbf{q})$ [5] as $\mathbf{q}=\operatorname*{arg\,min}_{\mathbf{q}}E(\mathbf{q})$, where $E(\mathbf{q})=E_{\mathrm{in}}(\mathbf{q})+E_{\mathrm{st}}(\mathbf{q})+E_{\mathrm{be}}(\mathbf{q})+E_{\mathrm{tw}}(\mathbf{q})$, and $E_{\mathrm{in}}(\mathbf{q})$, $E_{\mathrm{st}}(\mathbf{q})$, $E_{\mathrm{be}}(\mathbf{q})$, and $E_{\mathrm{tw}}(\mathbf{q})$ represent inertia, stretching, bending, and twisting objectives, respectively. We define these objectives in the following sections and provide their gradient in the supplementary material.
IV-A Inertia
We define the inertia objective as follows:
@@ IV-A Inertia
$E_{\mathrm{in}}(\mathbf{q})=\frac{1}{2\Delta t^{2}}\left\lVert\mathbf{q}-\mathbf{q}^{*}\right\rVert^{2}_{\mathbf{M}},$ (1)
@@
where $\mathbf{q}^{*}=\mathbf{q}^{t}+\Delta t\mathbf{\dot{q}}^{t}+\Delta t^{2}\mathbf{M}^{-1}\mathbf{f}_{\mathrm{ext}}$, $\mathbf{q}^{t}$ and $\mathbf{\dot{q}}^{t}$ denote the generalized positions and velocities at time $t$, respectively, $\Delta t$ the time step size, $\mathbf{M}$ a diagonal generalized mass matrix [7, 46], and $\mathbf{f}_{\mathrm{ext}}$ generalized external forces. While we set gravity as the only external force, one could specify spatially varying external loads and winds (even on edge angles) [5].
IV-B Stretching
We define the stretching objective [7] as $E_{\mathrm{st}}(\mathbf{q})=\sum_{i=1}^{N-2}E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})$, where $E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})$ is given by
$E_{\mathrm{st},i}(\mathbf{x}_{i},\mathbf{x}_{i+1})=\frac{1}{2}\left(\frac{s\boldsymbol{\alpha}_{i}\pi r^{2}}{\mathbf{\bar{l}}_{i}}\right)(\mathbf{l}_{i}-\mathbf{\bar{l}}_{i})^{2},$ (2)
@@ IV-B Stretching
with $s$ denoting a global constant scalar (defined in Sec. V-A) that scales stiffness parameters, $\boldsymbol{\alpha}_{i}$ as the stiffness parameter for stretching on edge $i$ (we use different stiffness parameters for stretching, bending, and twisting to enable finer control [47, 5], unlike [7] which uses only Young’s and shear modulus), $\mathbf{l}_{i}=\left\lVert\mathbf{x}_{i+1}-\mathbf{x}_{i}\right\rVert_{2}$ representing the length of edge $i$, and $\mathbf{\bar{l}}_{i}$ as its rest length.
IV-C Bending
@@
$E_{\mathrm{be},i}(\mathbf{y}_{i})=\frac{1}{2}\left(\frac{s\boldsymbol{\beta}_{i}\pi r^{4}}{4(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)\left\lVert\boldsymbol{\kappa}_{i}-\boldsymbol{\bar{\kappa}}_{i}\right\rVert_{2}^{2},$ (3)
and $\mathbf{y}_{i}=(\mathbf{x}_{i-1}^{T},\theta_{i-1},\mathbf{x}_{i}^{T},\theta_{i},\mathbf{x}_{i+1}^{T})^{T}\in\mathbb{R}^{11}$, $\boldsymbol{\beta}_{i}$ denotes the bending stiffness at vertex $i$, and $\boldsymbol{\kappa}_{i}$ and $\boldsymbol{\bar{\kappa}}_{i}$ are the four-dimensional curvature and rest curvature, respectively [6, 9]. We use the formulation with four-dimensional curvatures and rest curvatures [6, 9] because a simplified bending formulation with two-dimensional curvatures and rest curvatures via averaged material frames [7] is known to fail to evaluate strand bending correctly due to the non-unit and non-orthogonal averaged material frames [48, 12]. We provide more details on these bending formulations in the supplementary material.
IV-D Twisting